Why is the K.E acquired by a mass m travelling a certain distance d starting from rest under the action of constant force is independent of mass m?

Grade:11

1 Answers

Arun
6 years ago
Dear Ishan
 
a = F/m
 
Now
V² = u² +2*a*s
V = \sqrt(0 + 2*(F/m) *d
V = \sqrt(2Fd/m)
 
K.E. acquired will be
K.E. = (1/2) * m* v²
= (1/2) * m* (2Fd/m)
 = F*d
 
Independent of mass m
 
Regards
Arun (askIITians forum expert)